### v1

# Joint Transmission and Compression Design for 6G Networked Sensing with Limited-Capacity Backhaul

Abstract
This paper considers networked sensing in cellular network, where multiple base stations (BSs) first compress their received echo signals from multiple targets and then forward the quantized signals to the cloud via limited-capacity backhaul links, such that the cloud can leverage all useful echo signals to perform high-resolution localization. Under this setup, we manage to characterize the posterior Cramér-Rao Bound (PCRB) for localizing all the targets as a function of the transmit covariance matrix and the compression noise covariance matrix of each BS. Then, a PCRB minimization problem subject to the transmit power constraints and the backhaul capacity constraints is formulated to jointly design the BSs’ transmission and compression strategies. We propose an efficient algorithm to solve this problem based on the alternating optimization technique. Specifically, it is shown that when either the transmit covariance matrices or the compression noise covariance matrices are fixed, the successive convex approximation technique can be leveraged to optimize the other type of covariance matrices locally. Numerical results are provided to verify the effectiveness of our proposed algorithm.
I Introduction
Integrated sensing and communication (ISAC) is identified as one of the six key usage scenarios of the six-generation (6G) cellular systems [1]. The main challenge in ISAC is how to embed the sensing functionality into the future 6G communication systems efficiently [2, 3]. In the literature, plenty of works have studied the practical strategies to perform sensing[4, 5, 6] and the fundamental limits of sensing [7, 8, 9, 10, 11] in communication systems. It is worth noting that in most of the existing ISAC works, the sensing functionality is achieved via one base station (BS), just as in radar systems.
However, there exists a notable difference between the BSs and the conventional radars - the BSs across large-scale area are inter-connected via the backhaul links and can share information with each other. Actually, this property has been utilized to achieve cooperative communication in cellular network, giving rise to cloud radio access network (C-RAN), cell-free massive multiple-input multiple-output (MIMO), etc., where multiple BSs jointly serve the users via sharing the data messages and the channel state information (CSI). Similarly, in the future 6G ISAC systems, multiple BSs may share the local sensing information with each other to perform networked sensing [12, 13]. This is a major advantage of 6G sensing systems over the conventional radar systems.
In this paper, we consider a 6G networked sensing system consisting of multiple BSs and multiple targets, as shown in Fig. 1. Each BS first emits wireless signals in the downlink and then receives the echo signals from the targets. Next, each BS adopts a compress-then-forward strategy, under which it first compresses its received echo signals and then forwards the compression bits to the cloud via a limited-capacity backhaul link. At last, the cloud may utilize the global information observed by all the BSs to jointly localize the targets. Under this system, we manage to characterize the Posterior Cramér-Rao Bound (PCRB) [14] for localization as a function of the BSs’ transmit covariance matrices and compression noise covariance matrices. Based on this result, we formulate a PCRB minimization problem subject to each BS’s transmit power constraint and backhaul capacity constraint. Moreover, we propose an efficient algorithm to solve this problem via the alternating optimization technique. Numerical results are provided to verify the effectiveness of our proposed algorithm. In the literature, most of the ISAC works focus on the sensing limits in a single-BS and single-target scenario [7, 8, 9, 10, 11], where sensing is performed locally by the BS. Our work makes an early attempt to consider PCRB characterization under multi-BS and multi-target scenario and investigate the impact of limited-capacity backhaul on sensing design.
II System Model
Figure 1: System model for networked sensing with limited-capacity backhaul.
We consider the sensing job in a multi-cell ISAC system as shown in Fig. 1, where $N$ BSs, each equipped with $M_{t}$ transmit antennas and $M_{r}$ receive antennas, are connected to the cloud via the backhaul links and cooperatively localize $K$ targets via utilizing the wireless signals. We focus on a two-dimensional (2D) coordinate system, while the known position of BS $n$ is denoted as $\bm{\rho}_{n}=[\rho_{n}^{x},\rho_{n}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, and the unknown position of target $k$ is denoted as $\bm{q}_{k}=[q_{k}^{x},q_{k}^{y}]^{T}\in\mathbb{R}^{2\times 1}$, $n=1,\ldots,N$ and $k=1,\ldots,K$. In this paper, we assume that the unknown target positions $\bm{q}=[\bm{q}_{1}^{T},\dots,\bm{q}_{K}^{T}]^{T}\in\mathbb{R}^{2K\times 1}$ are random with a joint probability density function (PDF) of $p_{\bm{q}}(\bm{q})$. Moreover, the angle-of-departure (AOD) of the signal transmitted from BS $n$ to target $k$ and the angle-of-arrival (AOA) of the signal reflected from target $k$ to BS $n$ are denoted as
$\theta_{n,k}=\arctan\frac{q_{k}^{x}-\rho_{n}^{x}}{q_{k}^{y}-\rho_{n}^{y}},\quad\forall n,k.$ (1)
Moreover, define $\bm{v}(\theta_{n,k})\in\mathbb{C}^{M_{t}\times 1}$ and $\bm{a}(\theta_{n,k})\in\mathbb{C}^{M_{r}\times 1}$ as the transmit and receive steering vectors of BS $n$ towards target $k$, $\forall n,k$.
Under the above system, all the BSs transmit wireless signals simultaneously in the downlink, then receive the echo signals from the targets, and forward their received echo signals to the cloud via their backhaul links for performing localization. Specifically, let $\bm{x}_{n}\in\mathbb{C}^{M_{t}\times 1}$ with the covariance matrix $\bm{R}_{n}=\mathbb{E}[\bm{x}_{n}\bm{x}_{n}^{H}]$ denote the signal emitted by the transmit antenna array of BS $n$, $\forall n$. Then, the echo signals at the receive antenna array of BS $n$ can be written as
$\bm{y}_{n}\!=\!\bm{A}_{n}\sum_{u=1}^{N}\bm{B}_{n,u}\bm{V}_{u}^{T}\bm{x}_{u}+\bm{z}_{n}\!=\!\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}+\bm{z}_{n},\forall n.$ (2)
In the above, $\bm{V}_{n}=[\bm{v}(\theta_{n,1}),\ldots,\bm{v}(\theta_{n,K})]\in\mathbb{C}^{M_{t}\times K}$ and $\bm{A}_{n}=[\bm{a}(\theta_{n,1}),\ldots,\bm{a}(\theta_{n,K})]\in\mathbb{C}^{M_{r}\times K}$ denote the collection of the transmit and receive steering vectors of BS $n$, $\bm{B}_{n,u}=\textrm{diag}(\bm{b}_{n,u})\in\mathbb{C}^{K\times K}$, where $\bm{b}_{n,u}=[b_{n,u,1},\ldots,b_{n,u,K}]^{T}$ with $b_{n,u,k}$ being the attenuation coefficient from the transmit antenna array of BS $u$ to target $k$ to the receive antenna array of BS $n$, which contains both the round-trip path loss and the radar cross-section (RCS) of the target, $\bm{z}_{n}\sim\mathcal{CN}(0,\sigma^{2}\bm{I})$ is the Gaussian noise at BS $n$, and $\bm{G}_{n,u}\triangleq\bm{A}_{n}\bm{B}_{n,u}\bm{V}_{u}^{T}$, $\forall n,u$.
Due to the limited backhaul capacity, BS $n$ needs to compress the echo signal $\bm{y}_{n}$, $\forall n$. In this paper, we consider the Gaussian test channel to model the compression process [15]. Specifically, the quantized signal of BS $n$ is given as [15]
$\tilde{\bm{y}}_{n}$ $=\bm{y}_{n}+\bm{e}_{n}=\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}+\bm{z}_{n}+\bm{e}_{n},\forall n,$ (3)
where $\bm{e}_{n}\in\mathbb{C}^{M_{r}\times 1}$ is the Gaussian quantization error vector with zero mean and covariance matrix $\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, i.e., $\bm{e}_{n}\sim\mathcal{CN}(\bm{0},\bm{Q}_{n})$. Based on (3), the rate (in terms of bits/second/sample) for BS $n$ to transmit $\tilde{\bm{y}}_{n}$ via backhaul is
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{Q}_{n})=$
$\mathbb{E}_{\bm{q}}\left[\log_{2}\left(\frac{\left|\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}+\bm{Q}_{n}\right|}{|\bm{Q}_{n}|}\right)\right],~{}\forall n,$ (4)
where the expectation is over the target location vector $\bm{q}=[\bm{q}_{1}^{T},\dots,\bm{q}_{K}^{T}]^{T}$, which determines the channel $\bm{G}_{n,u}$’s.
After receiving the quantized signals from all the $N$ BSs, the cloud aims to localize all the $K$ targets. In the following, we first characterize the localization performance under networked sensing as a function of $\bm{R}_{1},\ldots,\bm{R}_{N}$ and $\bm{Q}_{1},\ldots,\bm{Q}_{N}$, and then optimize these covariance matrices to improve the sensing performance subject to the backhaul capacity constraints.
III Posterior Cramér-Rao Bound for Localization
In this work, we utilize the PCRB [14] as the performance metric to evaluate the performance of the networked sensing system described in the above. Specifically, let $\bm{q}=[\bm{q}_{1}^{T},\ldots,\bm{q}_{K}^{T}]^{T}$ denote the collection of target locations, and $\bm{b}^{\rm R}=[(\bm{b}^{\rm R}_{1,1})^{T},\ldots,(\bm{b}^{\rm R}_{1,N})^{T},\ldots,(\bm{b}^{\rm R}_{N,1})^{T},\ldots,(\bm{b}^{\rm R}_{N,N})^{T}]^{T}$ and $\bm{b}^{\rm I}=[(\bm{b}^{\rm I}_{1,1})^{T},\ldots,(\bm{b}^{\rm I}_{1,N})^{T},\ldots,(\bm{b}^{\rm I}_{N,1})^{T},\ldots,(\bm{b}^{\rm I}_{N,N})^{T}]^{T}$ denote the collection of the real parts and that of the imaginary parts of the attenuation coefficients, where $\bm{b}^{\rm R}_{n,u}=[b_{n,,1}^{\rm R},\ldots,b_{n,u,K}^{\rm R}]^{T}$ and $\bm{b}^{\rm I}_{n,u}=[b_{n,u,1}^{\rm I},\ldots,b_{n,u,K}^{\rm I}]^{T}$ with $b_{n,u,k}^{\rm R}$ and $b_{n,u,k}^{\rm I}$ being the real part and the imaginary part of $b_{n,u,k}$, $\forall n,u,k$. Moreover, define $\boldsymbol{\xi}=[\bm{q}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$ as the collection of all the unknown parameters to be estimated, and $\tilde{\bm{y}}=[\tilde{\bm{y}}_{1}^{T},\ldots,\tilde{\bm{y}}_{N}^{T}]^{T}$ as the collection of the de-quantized signals at the cloud, where $\tilde{\bm{y}}_{n}$ is given in (3). We assume that the target locations are random variables with a joint probability density function (PDF) $p_{\bm{q}}(\bm{q})$, and the attenuation coefficients are deterministic but unknown variables. In the rest of this section, we aim to characterize the PCRB for estimating $\boldsymbol{\xi}$ based on the de-quantized signals at the cloud $\tilde{\bm{y}}$ and the prior knowledge of PDF $p_{\bm{q}}(\bm{q})$.
In order to derive the PCRB for $\boldsymbol{\xi}$, we first calculate the Fisher information matrix (FIM) for estimating $\boldsymbol{\xi}$
$\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})=\bm{F}_{0,\boldsymbol{\xi}}+\bm{F}_{p},$ (5)
where $\bm{F}_{0,\boldsymbol{\xi}}$ denotes the FIM corresponding to the de-quantized signals given in (3)
$\bm{F}_{0,\boldsymbol{\xi}}=\mathbb{E}_{\tilde{\bm{y}},\bm{q}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\xi})}{\partial\boldsymbol{\xi}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\xi})}{\partial\boldsymbol{\xi}}\Bigg{)}^{T}\right],$ (6)
with
$p(\tilde{\bm{y}}|\boldsymbol{\xi})=$ $~{}\prod_{n=1}^{N}\frac{1}{\pi^{L}|\sigma^{2}\bm{I}+\bm{Q}_{n}|}\exp\Bigg{[}-\Bigg{(}\tilde{\bm{y}}_{n}-\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}\Bigg{)}^{H}$
×$\times(\sigma^{2}\bm{I}+\bm{Q}_{n})^{-1}\Bigg{(}\tilde{\bm{y}}_{n}-\sum_{u=1}^{N}\bm{G}_{n,u}\bm{x}_{u}\Bigg{)}\Bigg{]},$ (7)
and $\bm{F}_{p}$ denotes the FIM corresponding to the prior knowledge about location distribution
$\bm{F}_{p}$ $=\mathbb{E}_{\bm{q}}\left[\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\boldsymbol{\xi}}\Bigg{(}\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\boldsymbol{\xi}}\Bigg{)}^{T}\right]$
$=\begin{bmatrix}\mathbb{E}_{\bm{q}}\Bigg{\{}\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\bm{q}}\Big{(}\frac{\partial\log p_{\bm{q}}(\bm{q})}{\partial\bm{q}}\Big{)}^{T}\Bigg{\}}&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}.$ (8)
Given any PDF $p_{\bm{q}}(\bm{q})$, $\bm{F}_{p}$ can be computed based on (III). Therefore, the remaining job to characterize $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is to quantify $\bm{F}_{0,\boldsymbol{\xi}}$ given in (6). It is observed from (2) that $\bm{G}_{n,u}$’s are functions of $\bm{\theta}=[\theta_{1,1},\ldots,\theta_{1,N},\ldots,\theta_{N,1},\ldots,\theta_{N,K}]^{T}$. Moreover, because $\bm{\theta}$ is a complicated function of the target locations $\bm{q}$ as shown in (1), $p(\tilde{\bm{y}}|\boldsymbol{\xi})$ is a complicated function of $\bm{q}$, and it is challenging to calculate its derivative over $\bm{q}$. We apply the chain rule to tackle this challenge. Specifically, define $\boldsymbol{\zeta}=[\boldsymbol{\theta}^{T},(\bm{b}^{\rm R})^{T},(\bm{b}^{\rm I})^{T}]^{T}$. Moreover, define
$\bm{u}_{n,k}=\left[\frac{\partial\theta_{n,k}}{\partial q_{k}^{x}},\frac{\partial\theta_{n,k}}{\partial q_{k}^{y}}\right]^{T},~{}\forall n,k,$ (9)
where according to (1),
$\frac{\partial\theta_{n,k}}{\partial q_{n}^{x}}$ $=\frac{1}{(1+\tan^{2}\theta_{n,k})(q_{k}^{y}-\rho_{n}^{y})},$ (10)
$\frac{\partial\theta_{n,k}}{\partial q_{n}^{y}}$ $=-\frac{1}{(1+\cot^{2}\theta_{n,k})(q_{k}^{x}-\rho_{n}^{x})}.$ (11)
Then, based on the chain rule, $\bm{F}_{0,\boldsymbol{\xi}}$ can be obtained as follows
$\bm{F}_{0,\boldsymbol{\xi}}=\mathbb{E}_{\boldsymbol{\xi}}[\bm{U}\bm{F}_{0,\boldsymbol{\zeta}}\bm{U}^{T}].$ (12)
In the above, $\bm{U}$ is the Jacobian matrix given by
$\bm{U}\triangleq\frac{\partial\boldsymbol{\zeta}}{\partial\boldsymbol{\xi}}=\begin{bmatrix}\hat{\bm{U}}&\bm{0}\\ \bm{0}&\bm{I}\end{bmatrix}\in\mathbb{C}^{(2K+N^{2}K)\times(NK+N^{2}K)},$ (13)
where
$\hat{\bm{U}}=[\bm{U}_{1},\dots,\bm{U}_{N}]\in\mathbb{C}^{2K\times NK},$ (14)
with $\bm{U}_{n}$ being a block diagonal matrix whose $k$-th block diagonal component is $\bm{u}_{n,k}$, $\forall n,k$. Moreover, $\bm{F}_{0,\boldsymbol{\zeta}}$ is defined as
$\bm{F}_{0,\boldsymbol{\zeta}}=\mathbb{E}_{\tilde{\bm{y}}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}\Bigg{)}^{T}\right],$ (15)
where $p(\tilde{\bm{y}}|\boldsymbol{\zeta})$ has the similar form to $p(\tilde{\bm{y}}|\boldsymbol{\xi})$, but we should treat $\bm{G}_{n,u}$’s in (2) as functions of $\bm{\theta}$ instead of $\bm{q}$. The explicit expression for $\bm{F}_{0,\boldsymbol{\zeta}}$ can be given by the following proposition.
Proposition 1
Define $\bm{F}_{1}$, $\bm{F}_{2}$, and $\bm{F}_{3}$ as (16) and (17) on the top of the next page, where
$\bm{F}_{1}$ $=\begin{bmatrix}\bm{F}_{1}(1,1)&\dots&\bm{F}_{1}(1,N)\\ \vdots&\ddots&\vdots\\ \bm{F}_{1}(N,1)&\dots&\bm{F}_{1}(N,N)\end{bmatrix},\bm{F}_{2}=\begin{bmatrix}\bm{F}_{2}(1,1,1)&\dots&\bm{F}_{2}(1,1,N)&\dots&\bm{F}_{2}(1,N,1)&\dots&\bm{F}_{2}(1,N,N)\\ \vdots&\ddots&\vdots&\dots&\vdots&\ddots&\vdots\\ \bm{F}_{2}(N,1,1)&\dots&\bm{F}_{2}(N,1,N)&\dots&\bm{F}_{2}(N,N,1)&\dots&\bm{F}_{2}(N,N,N)\end{bmatrix},$ (16)
$\bm{F}_{3}$ $=\begin{bmatrix}\bm{F}_{3}(1,1,1,1)&\dots&\bm{F}_{3}(1,1,1,N)&\dots&\bm{F}_{3}(1,1,N,1)&\dots&\bm{F}_{3}(1,1,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(1,N,1,1)&\dots&\bm{F}_{3}(1,N,1,N)&\dots&\bm{F}_{3}(1,N,N,1)&\dots&\bm{F}_{3}(1,N,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(N,1,1,1)&\dots&\bm{F}_{3}(N,1,1,N)&\dots&\bm{F}_{3}(N,1,N,1)&\dots&\bm{F}_{3}(N,1,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(N,N,1,1)&\dots&\bm{F}_{3}(N,N,1,N)&\dots&\bm{F}_{3}(N,N,N,1)&\dots&\bm{F}_{3}(N,N,N,N)\end{bmatrix}.$ (17)
$\bm{F}_{1}(i,i)=\sum_{u=1}^{N}\Big{[}\Big{(}\bm{A}_{u}^{H}\bm{O}^{-1}_{u}\bm{A}_{u})\odot(\bm{B}^{*}_{ui}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{ui}\Big{)}\Big{]}$
$+(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{ii}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{ii})$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}^{*}_{ii}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{ii})$
$+\sum_{u=1}^{N}\Big{[}\Big{(}\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}_{iu}^{*}\bm{V}_{u}^{H}\bm{R}_{u}^{*}\bm{V}_{u}\bm{B}_{iu}\Big{)}\Big{]},\forall i,$ (18)
$\bm{F}_{1}(i,n)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{in}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\dot{\bm{V}}_{n}\bm{B}_{in})$
$+(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\dot{\bm{A}}_{n})\odot(\bm{B}^{*}_{ni}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{ni}),\forall i\neq n,$ (19)
$\bm{F}_{2}(i,i,i)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{ii}^{*}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i})$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{ii}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i,$ (20)
$\bm{F}_{2}(i,i,n)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{in}^{*}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\bm{V}_{n}),\forall i\neq n,$ (21)
$\bm{F}_{2}(i,n,i)=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{B}_{ni}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i\neq n,$ (22)
$\bm{F}_{2}(i,n,k)=\mathbf{0},\forall i\neq n,i\neq k,$ (23)
$\bm{F}_{3}(n,i,n,i)=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i,$ (24)
$\bm{F}_{3}(n,i,n,k)=\mathbf{0},\forall i\neq k,$ (25)
$\bm{F}_{3}(n,i,u,k)=\mathbf{0},\forall n\neq u,$ (26)
with $\dot{\bm{A}}_{n}=[\frac{\partial\bm{a}(\theta_{n,1})}{\partial\theta_{n,1}},\ldots,\frac{\partial\bm{a}(\theta_{n,K})}{\partial\theta_{n,K}}]\in\mathbb{C}^{M_{r}\times K}$ and $\bm{O}_{n}=\sigma^{2}\bm{I}+\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, $\forall n$. Then, $\bm{F}_{0,\boldsymbol{\zeta}}$ given in (15) can be expressed as
$\bm{F}_{0,\boldsymbol{\zeta}}=2\begin{bmatrix}\Re\{\bm{F}_{1}\}&\Re\{\bm{F}_{2}\}&-\Im\{\bm{F}_{2}\}\\ \Re\{\bm{F}_{2}\}^{T}&\Re\{\bm{F}_{3}\}&-\Im\{\bm{F}_{3}\}\\ -\Im\{\bm{F}_{2}\}^{T}&-\Im\{\bm{F}_{3}\}^{T}&\Re\{\bm{F}_{3}\}\end{bmatrix},$ (27)
where given each matrix $\bm{X}$, $\Re\{\bm{X}\}$ and $\Im\{\bm{X}\}$ are the matrices with each element being the real part and imaginary part of the element in $\bm{X}$, respectively.
Proof:
The details of proof are omitted due to limited pages. ∎
After characterizing the FIM $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ shown in (5), where $\bm{F}_{0,\boldsymbol{\xi}}$ is obtained via Proposition 1 and $\bm{F}_{p}$ is obtained via (III), the PCRB for the mean-squared error (MSE) of any estimator of $\bm{q}$, denoted by $\hat{\bm{q}}$, can be expressed as
$\mathbb{E}_{\bm{q}}[\|\hat{\bm{q}}-\bm{q}\|^{2}]\geq\sum_{i=1}^{2K}[(\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N}))^{-1}](i,i),$ (28)
where for any squared matrix $\bm{X}$, $\bm{X}(i,i)$ denotes its $i$-th diagonal element.
Remark 1
We want to emphasize that our result is a generalization of the result in [16]. Specifically, the work [16] has characterized the Cramér-Rao Bound for estimating the unknown but deterministic locations of targets for the special case of one BS, i.e., $N=1$, while our result holds for estimating the random target locations in the general case of networked sensing via cooperation among multiple BSs.
IV Optimizing Transmission and Compression Strategies for PCRB Minimization
In this section, we aim to design the transmit covariance matrices and the compression noise covariance matrices of all the BSs, i.e., $\bm{R}_{n}$ and $\bm{Q}_{n}$, $\forall n$, to minimize the PCRB for localization, subject to the individual transmit power constraint and backhaul capacity constraint of each BS. Specifically, the above problem can be formulated as
[…]
s.t. $~{}\text{tr}(\bm{R}_{n})\leq\bar{P}_{n},~{}\forall n,$ (29b)
$~{}D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{Q}_{n})\leq\bar{D}_{n},~{}\forall n,$ (29c)
$~{}\bm{Q}_{n}\succeq\bm{0},~{}\forall n,$ (29d)
where $\bar{P}_{n}$ is transmit power constraint for BS $n$, $\bar{D}_{n}$ is the backhaul capacity constraint for BS $n$, and $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is the backhaul transmission rate of BS $n$ given in (II), $\forall n$.
Note that the objective function of Problem (P1) is a complicated function of $\bm{R}_{n}$’s and $\bm{Q}_{n}$’s. To tackle this challenge, we adopt the Schur complement technique to equivalently transform Problem (P1) to the following problem
[…]
s.t. $~{}\begin{bmatrix}\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (30b)
[…]
where $t_{1},\ldots,t_{2K}$ are auxiliary variables and $\bm{e}_{i}$ is the $i$th column of the identity matrix $\bm{I}$, $\forall i$. Problem (P2) is still challenging, because all the variables are coupled together. In this work, we design an alternating optimization based algorithm to iteratively solve Problem (P2). Specifically, at each iteration of the alternating optimization based algorithm, we first fix $\bm{Q}_{n}$’s and optimize $\bm{R}_{n}$’s and $t_{i}$’s in Problem (P2), and then fix $\bm{R}_{n}$’s and optimize $\bm{Q}_{n}$’s and $t_{i}$’s in Problem (P2). This can guarantee that the objective value of Problem (P2) will be decreased after each iteration. The algorithm will terminate at some iteration when the decrease in the objective value is smaller than some threshold. In the following, we respectively show how to optimize $\bm{R}_{n}$’s and $t_{i}$’s given $\bm{Q}_{n}$’s and how to optimize $\bm{Q}_{n}$’s and $t_{i}$’s given $\bm{R}_{n}$’s in Problem (P2).
IV-A Optimization of Transmit Covariance Matrices
Given any $\bm{Q}_{n}=\bar{\bm{Q}}_{n}$’s, we aim to solve the following sub-problem of Problem (P2)
[…]
Problem (P2.1) is a non-convex problem because $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$’s given in (II) are concave functions, instead of convex functions, of $\bm{R}_{n}$’s. We apply the successive convex approximation (SCA) technique to tackle this challenge. Specifically, because $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$ is a concave function, given any $\tilde{\bm{R}}_{n}\succeq\bm{0}$, $\forall n$, the following inequality holds
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})\leq\hat{D}_{n}(\{\bm{R}_{n},\tilde{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n}),~{}\forall n,$ (32)
where
$\hat{D}_{n}(\{\bm{R}_{n},\tilde{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})=$
$\mathbb{E}_{\bm{q}}\Bigg{\{}\text{tr}\Bigg{[}\bar{\bm{\Sigma}}_{n}^{-1}\Bigg{(}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}+\bar{\bm{Q}}_{n}\Bigg{)}\Bigg{]}\Bigg{\}}/\log 2$
$-M_{r}/\log 2+\mathbb{E}_{\bm{q}}\Big{[}\log_{2}|\bar{\bm{\Sigma}}_{n}|\Big{]}-\log_{2}|\bar{\bm{Q}}_{n}|,\quad\forall n,$ (33)
$\bar{\bm{\Sigma}}_{n}=\sum\limits_{u=1}^{N}\bm{G}_{n,u}\tilde{\bm{R}}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}_{n}+\bar{\bm{Q}}_{n}.$ (34)
Based on the above inequality, given any $\tilde{\bm{R}}_{n}\succeq\bm{0}$, $\forall n$, we aim to solve the following relaxed problem of Problem (P2.1)
$\min_{\{\bm{R}_{n}\}_{n=1}^{N},\{t_{i}\}_{i=1}^{2K}}~{}$ $\sum_{i=1}^{2K}t_{i}$ (35a)
s.t. $\hat{D}_{n}(\{\bm{R}_{n},\tilde{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})\leq\bar{D}_{n},\forall n,$ (35b)
[…]
Note that in Problem (35), the non-convex constraint (29c) in Problem (P2.1) is replaced by the convex constraint (35b). Because Problem (35) is a convex problem, we can utilize the interior-point method to globally solve it.
We summarize the SCA-based algorithm to solve Problem (P2.1) via Problem (35) in Algorithm I. According to the theory about the SCA technique, Algorithm I can yield a locally optimal solution to Problem (P2.1).
TABLE I: Algorithm I: SCA-Based Algorithm for Problem (P2.1)
  • 1.
    Initialize: Set $\bm{R}^{(0)}_{n}$ as the initial matrix of $\bm{R}_{n}$, $\forall n$, and $m=0$, where $m$ denotes the index of iteration;
  • 2.
    Repeat
    • a.
      Set $m=m+1$;
    • b.
      Update $\tilde{\bm{R}}_{n}=\bm{R}^{(m-1)}_{n}$, $\forall n$, in Problem (35);
    • c.
      Update $\bm{R}^{(m)}_{1},\ldots,\bm{R}^{(m)}_{N}$ as the optimal solution of Problem (35);
  • 3.
    Until convergence.
IV-B Optimization of Compression Noise Covariance Matrices
Next, we show how to optimize the compression noise covariance matrices. Given any $\bm{R}_{n}=\bar{\bm{R}}_{n}$’s, we aim to solve the following sub-problem of Problem (P2)
[…]
Problem (P2.2) is non-convex because $\bm{F}_{\boldsymbol{\xi}}(\{\bar{\bm{R}}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is not a concave function of $\bm{Q}_{n}$’s in (30b), and $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is not a convex function of $\bm{Q}_{n}$ in (29c). To tackle the above challenge, we first change the optimization variables. Specifically, define
$\bm{T}_{n}=(\sigma^{2}\bm{I}+\bm{Q}_{n})^{-1},~{}\forall n.$ (37)
By replacing $\bm{Q}_{n}$ with $\bm{T}_{n}^{-1}-\sigma^{2}\bm{I}$, $\forall n$, in $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ that is characterized in Proposition 1, we can get $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{T}_{n}\}_{n=1}^{N})$. Moreover, the backhaul compression rate given in (II) becomes
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{T}_{n})=\log_{2}\Bigg{|}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\bm{T}_{n}^{-1}\Bigg{|}$
$-\log_{2}\Big{|}\bm{T}_{n}^{-1}-\sigma^{2}\bm{I}\Big{|},~{}\forall n.$ (38)
Therefore, Problem (P2.2) is equivalent to
$\min_{\{\bm{T}_{n}\}_{n=1}^{N},\{t_{i}\}_{i=1}^{2K}}$ $~{}\sum_{i=1}^{2K}t_{i}$ (39a)
s.t. $\begin{bmatrix}\bm{F}_{\boldsymbol{\xi}}(\{\bar{\bm{R}}_{n},\bm{T}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (39b)
$D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{T}_{n})\leq\bar{D}_{n},~{}\forall n,$ (39c)
$\sigma^{-2}\bm{I}-\bm{T}_{n}\succeq\bm{0},\forall n.$ (39d)
Now, (39b) is convex over $\bm{T}_{n}$’s. However, (39c) is still non-convex because $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{T}_{n})$ is a concave function, instead of convex function, of $\bm{T}_{n}$. Similar to the optimization of $\bm{R}_{n}$’s, we apply the SCA technique to tackle this challenge. Specifically, the relaxed problem of Problem (39) is
$\min_{\{\bm{T}_{n}\}_{n=1}^{N},\{t_{i}\}_{i=1}^{2K}}$ $~{}\sum_{i=1}^{2K}t_{i}$ (40a)
s.t. $\tilde{D}_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{T}}_{n},\bm{T}_{n})\leq\bar{D}_{n},~{}\forall n,$ (40b)
[…]
where
$~{}\tilde{D}_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{T}}_{n},\bm{T}_{n})=$
$\mathbb{E}_{\bm{q}}\Big{\{}\text{tr}\left[\bar{\bm{\Omega}}_{n}^{-1}\Big{(}\bm{I}+\bar{\bm{J}}_{n}^{\frac{1}{2}}\bm{T}_{n}\bar{\bm{J}}_{n}^{\frac{1}{2}}\Big{)}\right]\Big{\}}/\log 2-M_{r}/\log 2$
$+\mathbb{E}_{\bm{q}}\Big{[}\log_{2}|\bar{\bm{\Omega}}_{n}|\Big{]}-\log_{2}\left|\bm{I}-\sigma^{2}\bm{T}_{n}\right|,~{}\forall n,$ (41)
$~{}\bar{\bm{J}}_{n}=\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bar{\bm{R}}_{u}\bm{G}_{n,u}^{H},~{}\forall n,$ (42)
$~{}\bar{\bm{\Omega}}_{n}=\bm{I}+\bar{\bm{J}}_{n}^{\frac{1}{2}}\bar{\bm{T}}_{n}\bar{\bm{J}}_{n}^{\frac{1}{2}},~{}\forall n.$ (43)
Problem (40) is a convex problem and can be solved by interior-point method. Then, the SCA-based algorithm to solve Problem (39) is summarized in Algorithm II. After the solution to Problem (39) is obtained, the compression noise covariance matrices can be obtained via (37).
TABLE II: Algorithm II: SCA-Based Algorithm for Problem (P2.2)
  • 1.
    Initialize: Set $\bm{T}^{(0)}_{n}$ as the initial matrix of $\bm{T}_{n}$, $\forall n$, and $m=0$, where $m$ denotes the index of iteration;
  • 2.
    Repeat
    • a.
      $m=m+1$;
    • b.
      Update $\tilde{\bm{T}}_{n}=\bm{T}^{(m-1)}_{n}$, $\forall n$, in Problem (40);
    • c.
      Update $\bm{T}^{(m)}_{1},\ldots,\bm{T}^{(m)}_{N}$ as the optimal solution of Problem (40);
  • 3.
    Until convergence. Set the compression noise covariance matrices as $\bm{Q}_{n}=(\bm{T}^{(m)}_{n})^{-1}-\sigma^{2}\bm{I}$, $\forall n$.
Finally, the alternating optimization based algorithm to solve Problem (P2) is summarized in Algorithm III.
TABLE III: Algorithm III: Overall Algorithm for Problem (P2)
  • 1.
    Initialize: Set $\bm{R}_{n}^{(0)}$ and $\bm{Q}_{n}^{(0)}$ as the initial matrices of $\bm{R}_{n}$ and $\bm{Q}_{n}$, $\forall n$, and $m=0$, where $m$ denotes the index of iteration;
  • 2.
    Repeat
    • a.
      $m=m+1$;
    • b.
      Set $\bar{\bm{Q}}_{n}=\bm{Q}_{n}^{(m-1)}$, $\forall n$, and update $\bm{R}^{(m)}_{n}$, $\forall n$, as the solution to Problem (P2.1) obtained via Algorithm I;
    • c.
      Set $\bar{\bm{R}}_{n}=\bm{R}_{n}^{(m)}$, $\forall n$, and update $\bm{Q}^{(m)}_{n}$, $\forall n$, as the solution to Problem (P2.2) obtained via Algorithm II;
  • 3.
    Until convergence.
V Numerical Results
In this section, we provide numerical examples to verify our results. The system consists of $N=2$ BSs, each equipped with $M_{t}=8$ transmit antennas and $M_{r}=8$ receive antennas, and $K=2$ targets. The locations of these two BSs are $\big{(}\frac{\sqrt{3}}{2},0\big{)}$ km and $\big{(}-\frac{\sqrt{3}}{2},0\big{)}$ km, respectively, and the two targets are randomly located at a small area around $\big{(}\frac{\sqrt{3}}{4},\frac{3}{4}\big{)}$ km and a small area around $\big{(}0,\frac{3}{4}\big{)}$ km, respectively. The power spectrum of the noise at the BSs is $-169$ dBm/Hz, and the channel bandwidth is $1$ MHz. We assume that the two BSs possess the same transmit power constraint, denoted by $\bar{P}=\bar{P}_{1}=\bar{P}_{2}$, and the same backhaul capacity constraint, denoted by $\bar{D}=\bar{D}_{1}=\bar{D}_{2}$.
To verify the effectiveness of Algorithm III to solve Problem (P2), we provide three benchmark schemes. Under Benchmark Scheme I, we perform uniform quantization to the received signals of each BS receive antenna as shown in [17] and merely optimize the transmit covariance matrices based on Algorithm I. Under Benchmark Scheme II, we fix the transmit covariance matrices of the BSs as $\bm{R}_{1}=\bm{R}_{2}=\frac{\bar{P}}{M_{t}}\bm{I}$ and merely optimize the compression noise covariance matrices based on Algorithm II. Under Benchmark Scheme III, we assume an ideal case where the backhaul capacity of each BS is infinite, i.e., constraint (29c) can be ignored in Problem (P2) and we merely design the transmit covariance matrices.
Figure 2: Sensing performance versus BSs’ transmit power.
Fig. 2 shows the PCRB performance achieved by various schemes versus BSs’ transmit power, when the common backhaul capacity is set as $\bar{D}=8$ bits per second per sample. It is observed that our proposed algorithm achieves much lower PCRB compared to Benchmark Schemes I and II under different transmit power, while almost the same PCRB as the ideal case achieved by Benchmark Scheme III.
Moreover, Fig. 3 shows the PCRB performance achieved by various schemes versus BSs’ backhaul capacity, when the common transmit power is set as $\bar{P}=31$ dBm. Similarly, it is observed that our proposed algorithm achieves much lower PCRB compared to Benchmark Schemes I and II when the backhaul capacity is moderate. When the backhaul capacity is sufficiently large, our proposed algorithm and Benchmark Schemes I and III achieve similar PCRB because the optimization of compression noise covariance matrices brings little gain in this case.
Figure 3: Sensing performance versus BSs’ backhaul capacity.
VI Conclusions
This work considered networked sensing in 6G cellular network, where multiple BSs forward their received echo signals to the cloud via limited-capacity backhaul links, and the cloud localizes all the targets based on the global information about the echo signals received by all the BSs. We characterized the PCRB for the sensing performance in this setup, and designed an efficient algorithm to optimize the BS transmit covariance matrices and compression noise covariance matrices for minimizing this PCRB subject to backhaul capacity constraints. Numerical results were provided to verify the effectiveness of our proposed scheme.